a beam of monochromatic light from a laser illuminates a piece of glass . the piece of glass is L= 25.21×10^-2m long with an index of refraction of n=2.83 . how long will a short pulse of the light beam take to travel the length of the piece of glass.

Express your answer in ns (1 ns = 1×10^-9s) and to two decimal places .

so if your calculator shows you 1.234×10^-9 multiply your answer by 1×10^-9 and enter that , 1.23 in the box below
Answer #1

Speed of light pulse in glass = 3*10^8/2.83

Time = distance /speed =25.21*10^-2/ 3*10^8/2.83

Time = 2.378*10^-9 s